OpenStudy (anonymous):

solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (anonymous):

could somrbody please help me on this question:)

OpenStudy (mathstudent55):

For this system, the elimination method works well. We need to add the equations and eliminate a variable. Since y + y = 2y, that does bnot eliminate y, but if we multiply an equation by -1, then we'll have y + (-y) = 0 which eliminates the y, Then we solve for x. First rewrite the first equation and write the second equation multiplied by -1 below it. -2x + y = 1 4x - y = 1 --------------(add) 2x = 2 Divide both sides by 2: x = 1 Now use substitution to replace x with 1 in the first original equation and solve for y: -2x + y = 1 -2(1) + y = 1 -2 + y = 1 y = 3 Solution: x = 1, y = 3